OpenStudy (osanseviero):

Can you calculate any solid of revolution both by shell and by disks? With example

OpenStudy (osanseviero):

I calculated the solid of revolution of 2 - x/2 between 1 and 2 with disks. How could I do this with shells?

OpenStudy (freckles):

of y=2-x/2?

OpenStudy (freckles):

between x=1 and x=2?

OpenStudy (freckles):

 and I guess you want to rotate the region bounded by y=2-x/2 and x=1 and x=2 and y=0 about the x-axis (or also known as y=0)